LBC CCD Cameras Maintenance Summer 2018

  • Date: August 18-22, 2018
  • Location: Electronic Lab Room 116
  • Performed by: Fernando Pedichini & Gustavo Rahmer.

Notes

  • Visually inspected the area around the detectors on both cameras for evidence of mylar foil residues. None found.
  • Checked proper operation of shutters.
  • No evidence found of unusual high noise (IT#6976) on LBCB or banding (IT#6748) on LBCR.
  • An old version of IDL was installed in the Lab PC, so the date of the computer had to be changed to 2015 to make the IDL license work. Therefore the FITS file names are off by 3 years (month and day are OK).

RED SCIENCE

Channel 1 2 3 4
Housekeeping OFF
Bias level (ADUs) 304 299 300 301
RON (ADUs) 4.9 4.8 5.2 4.7
Offset (V) 2.5877 2.6115 2.6150 2.5890
Housekeeping ON
Bias level (ADUs) 305 300 300 302
RON (ADUs) 5.1 4.8 5.0 4.7

Power Supplies (measured with oscilloscope)
Supply Voltage measured (V) RMS (V) @20Mhz
+8 VD 9.8 0.4
+5 VD 5.0 0.3
+11 VA 11.1 0.3
-11 VA -10.9 0.3
+25 VA 25.1 0.8
-25 VA -24.7 0.8

RED TECH

Housekeeping OFF
Channel 1 2
Bias level (ADUs) 301 301
RON (ADUs) 4.7 5.3
Offset (V) 2.5353 2.5525

BLUE SCIENCE

Channel 1 2 3 4
Housekeeping OFF
Bias level (ADUs) 298 296 299 298
RON (ADUs) 4.6 4.9 4.9 5.3
Offset (V) 2.5592 2.5620 2.5703 2.5640
Housekeeping ON
Bias level (ADUs) 297 297 303 299
RON (ADUs) 4.8 5.3 5.1 5.2

Power Supplies (measured with oscilloscope)
Supply Voltage measured (V) RMS (V) !@20Mhz
+8 VD 9.7 0.4
+5 VD 5.4 0.3
+11 VA 11.1 0.3
-11 VA -10.9 0.3
+25 VA 25.3 0.8
-25 VA -24.8 0.8

BLUE TECH

Housekeeping OFF
Channel 1 2
Bias level (ADUs) 300 299
RON (ADUs) 5.3 6.5
Offset (V) 2.5353 2.5724
--


%USERSIG{GustavoRahmer - 2018-08-22}%

Comments


 
Topic revision: r2 - 07 Aug 2019, GustavoRahmer
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback